OpenStudy (samirahdanyel):

The manager of a grocery store reports that there is a 12 percent chance that a customer buys apples during a shopping trip, a 5 percent chance that a customer buy apples and carrots, and a 17 percent chance that a customer buys apples or carrots. What is the probability of a customer buying carrots? percent

jimthompson5910 (jim_thompson5910):

hint: P(A and B) = P(A) + P(B) - P(A or B)

OpenStudy (samirahdanyel):

so it would be p (12 and 17) =P(12)+P(17)-P(12-17)

jimthompson5910 (jim_thompson5910):

A = event person buys apples B = event person buys carrots `The manager of a grocery store reports that there is a 12 percent chance that a customer buys apples during a shopping trip` means P(A) = 0.12 `...5 percent chance that a customer buy apples and carrots` so P(A and B) = 0.05 `...7 percent chance that a customer buys apples or carrots.` so P(A or B) = 0.07

jimthompson5910 (jim_thompson5910):

let x = P(B)

OpenStudy (samirahdanyel):

i got 0.10

jimthompson5910 (jim_thompson5910):

using the info above, we can go from P(A and B) = P(A) + P(B) - P(A or B) to 0.05 = 0.12 + x - 0.07

OpenStudy (samirahdanyel):

do we have to solve this equation

jimthompson5910 (jim_thompson5910):

my bad, I meant to say 0.17 instead of 0.07

jimthompson5910 (jim_thompson5910):

0.05 = 0.12 + x - 0.17 is the equation you need to solve

jimthompson5910 (jim_thompson5910):

solving 0.05 = 0.12 + x - 0.17 for x gives x = 0.10 so you are correct

OpenStudy (samirahdanyel):

i got 1/10

OpenStudy (samirahdanyel):

so its 10 percent?

jimthompson5910 (jim_thompson5910):

so 10%

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (samirahdanyel):

thank you for your help

jimthompson5910 (jim_thompson5910):

you're welcome